A vending machine contains nickels and dimes worth $14.50. There are 95 more nickels than dimes. How many nickels and how many dimes are there?
Answer by nerdybill(7384)

A vending machine contains nickels and dimes worth $14.50. There are 95 more nickels than dimes. How many nickels and how many dimes are there?
.
Let x = number of dimes
then
x+95 = number of nickels
.
.10x + .05(x+95) = 14.50

.10x + .05x+4.75 = 14.50
.15x+4.75 = 14.50
.15x = 9.75
x = 65 (number of dimes)
.
number of nickels:
x+95 = 65+95 = 160